QUESTION 5 1 pointsSave Answer What is the probabiity of getting a total of 6 or 9 when a pair of fair dice is tossed? 2 a.5/36 b. 20/36 О с. 4/36 d.9/36
Answer #1

We have total n=6^2= 36 possible outcomes when we toss 2 fair dice.

List the possible outcomes to get a total of 6 or 9 are{ (1,5),(2,4),(3,3),(4,2),(5,1),(3,6),(4,5),(5,4)(6,3)}

m=We have total 9 possible outcomes

So probability of getting total of 6 or 9 is =m/n

=9/ 36

Option d